### v2

@@ 1 Introduction
With the success of machine learning, (sub)gradient-type algorithms, such as (sub)gradient descent, coordinate descent, Frank-Wolf algorithm, and other variants, have gained much attention over the past decades [2, 3]. For smooth and/or convex cases, (sub)gradient methods are most well-understood. It is well-known that the basic gradient descent algorithm achieves linear convergence for minimizing smooth and strongly convex functions [4]. However, strong convexity is often not satisfied in practice, and many fundamental models in machine learning lack this good property [5].
@@
This motivates the study of different regularity conditions that are weaker than strong convexity, such as Polyak-Łojasiewicz inequality [6], restricted secant inequality [7], subdifferential error bound [8], and quadratic growth [9]. These regularity conditions go beyond strong convexity, and they mainly involve the relationship concerning the distance to the solution set, the cost value gap, and the length of the (sub)gradient. They often play a crucial step in many algorithm analyses. It has been shown that under those regularity conditions, a range of first-order optimization algorithms still enjoy linear convergence in many cases; see e.g., [6, 10, 11, 8, 12, 13, 14, 15]. Moreover, these convergence analyses even become more elegant under suitable regularity conditions; examples include the gradient descent method for smooth optimization [4, Theorem 2.1.15]; the proximal gradient method for nonsmooth convex optimization [8, Theorem 3.2]; and the proximal point method (PPM) to solve $0\in T(x)$ where $T$ is a maximal monotone set value mapping [16].
@@
Recent studies have revealed the relationship between some regularity conditions under various assumptions on the smoothness and convexity of the objective function [10, 8, 17, 18, 19, 20, 21, 22, 23]. We postpone the comparison between different studies after we formally define regularity conditions (see Remark 3.1). Different conditions are useful to prove the convergence of various algorithms. For example, the subdifferential error bound yields a clean proof for the proximal point method (PPM) [24, 16]; the step-length type error bound becomes useful for the proximal gradient method [8]; Polyak-Łojasiewicz inequality leads to a clean proof of the gradient descent [6, 10]; quadratic growth plays an important role in the analysis of proximal bundle methods [12] and spectral bundle methods [15, 14]. In some cases, it may be easier to establish the convergence proof for one regularity condition from another one, e.g. [25, Proposition 3.3]. Therefore, understanding the relationship or equivalence between common regularity conditions becomes crucial in algorithm analyses.
@@
Despite the results mentioned above on the equivalence [10, 8, 17, 18, 19, 20, 21, 22, 23], their scope remains limited to one of the following scenarios: 1) smooth functions, or 2) nonsmooth but convex functions. While smooth and convex (but not strongly convex) problems cover a variety of applications, practical applications routinely deal with problems that lack both qualities (e.g., training nonsmooth and nonconvex deep networks). Recent studies have further identified another amenable problem class: weakly convex functions [26]. This class of functions includes all convex functions, $L$-smooth functions, certain compositions of convex functions with smooth functions, and many cost functions in modern machine learning [27, 28]. Some algorithms for convex functions can also be extended to solve weakly convex functions, e.g., the classical bundle method [29], the stochastic projected subgradient method [30], and the damped proximal augmented Lagrangian method [31]. A recent study also proposes a unified analysis framework to analyze the convergence of algorithms designed for weakly convex functions [28]. In this paper, we aim to generalize the previous results on the equivalence of different regularity conditions [10, 8, 17, 18, 19, 20, 21] from the class of smooth functions (and the class of convex functions) to the class of nonsmooth weakly convex functions. These findings will benefit algorithm analyses from various perspectives.
@@
To demonstrate the equivalence between different regularity conditions, we revisit the classical (inexact) PPM, a conceptually simple algorithm that serves as a guideline for many algorithm developments [32]. In particular, we provide a simple new proof for the linear convergence in both the convex and weakly convex settings using Polyak-Łojasiewicz inequality, subdifferential error bound, and quadratic growth. Our proof is partially inspired by a recent result on spectral bundle methods [15, 14] and the analysis in [16]. As a by-product of our investigation, we clarify some subtleties in inexact PPM in terms of the inexactness and feasibility, which are less emphasized in the literature. In this paper, our contributions are as follows.
@@
  • We first clarify the relationship among common regularity conditions, including strongly convex, restricted secant inequality, subdifferential error bound, Polyak-Łojasiewicz inequality, and quadratic growth, in the class of weakly convex functions (Theorem 3.1). Most of the proofs follow from some elementary analysis. The connection between Polyak-Łojasiewicz inequality and subdifferential error bound utilizes the notion of slope in [21, Section 2] and requires some sophisticated arguments. We also derive the coefficient conversion among different conditions (see Table 1).
@@
  • Building upon the equivalence for different regularity conditions, we present a simple and modular proof for linear convergence of the PPM for convex optimization [24] (Theorem 4.2). The linear convergence also extends to the class of weakly convex functions with a proper initialization (Theorem 4.3). Our linear convergence result allows non-unique minimizers which is more general than a classical assumption of local Lipschitz continuity in [24] and also relaxes an upper Lipschitz condition in [33].
@@
  • Finally, we clarify some subtleties in the inexact PPM regarding inexactness stopping criteria and feasibility (Section 5.1). We further establish the linear convergence of inexact PPM (Theorem 5.3) under standard assumptions on the inexactness of the subproblem. Our proof is clean and modular, which combines a key inequality in [33] with Theorem 4.2.
@@
A preliminary summary of this work appeared in [1]. In this paper, we 1) complete the detailed proofs for all theorems and derive coefficient conversion among different conditions in Table 1, 2) extend the linear convergence result of the PPM from optimization on convex functions to weakly convex functions (Theorem 4.3), and 3) provide extra technical discussions on inexact PPM (Section 5.1). The rest of this paper is structured as follows. Section 2 summarizes the linear convergence of gradient descent as motivation and introduces some preliminaries in nonsmooth optimization techniques. Section 3 presents the relationship among different regularity conditions. Sections 4 and 5 focus on the (inexact) PPM and establish the sublinear and linear convergences. Section 6 presents three numerical experiments. Section 7 concludes this paper. Some auxiliary proofs and discussions are provided in Appendices D, A, D, C and B.
@@ 2.1 Motivation: Linear convergence of gradient descent algorithms
where $t_{k}>0$ is the step size. A textbook result says that when choosing a constant step size $t_{k}=\frac{1}{L}$, the GD iterates converge to $f^{\star}$ with a sublinear rate (precisely, $f(x_{k})-f^{\star}\leq L\mathrm{dist}^{2}(x_{0},S)/(2k)$); see e.g., [4, Corollary 2.1.2]. If the function $f$ is strongly convex, GD achieves a global linear convergence [4, Theorem 2.1.15].
However, the assumption of strong convexity is often not satisfied in practice. It is known that some alternative weaker assumptions are sufficient for linear convergences. We here introduce two notions: restricted secant inequality (RSI) [7], and Polyak-Łojasiewicz (PL) inequality [6]. A differentiable function $f$ satisfies RSI if there exists $\mu_{\mathrm{r}}>0$ such that
@@
Note that both RSI 2 and PL 3 imply that any stationary point of $f$ is a global minimum. However, they do not imply the uniqueness of stationary points or the convexity of the function. One can think that RSI 2 (resp. PL 3) requires that the gradient $\nabla f(x)$ grows faster than that of quadratic functions when moving away from the solution set $S$ (resp. the optimal value $f^{\star}$). Linear convergence of GD under the PL inequality was first proved in [6], and linear convergence under RSI was discussed in [34, Proposition 1] and [35, Proposition 1]. We summarize a simple version below.
@@
Thanks to RSI 2 and PL inequality 3, the proof of Theorem 2.1 is very elegant, taking only a few lines. For self-completeness, we provide a simple proof and some additional discussions in Section B.1. In particular, the RSI 2 leads to a quick proof of 4a, and the PL 3 allows for a simple proof of 4b. It is known that for $L$-smooth convex functions, the two conditions RSI 2 and PL 3 are equivalent (cf. [10, Theorem 2]). Some recent studies, such as [18, 9, 36], have explored the relationships among different regularity conditions for linear convergences. A nice summary appeared in [10, Theorem 2], but it only works in the context of $L$-smooth functions.
@@ 2.2 The notion of slope and Ekeland’s variational principle
We first introduce the Fréchet subdifferential. Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed function. We define the Fréchet subdifferential (see e.g., [37, Page 27]):
^$\hat{\partial}f(x):=\mathopen{}\mathclose{{}\left\{s\in\mathbb{R}^{n}\mid\liminf_{y\to x}\frac{f(y)-f(x)-\mathopen{}\mathclose{{}\left\langle s,y-x}\right\rangle}{\|y-x\|}\geq 0}\right\}.$
If $f$ is convex, the Fréchet subdifferential ^$\hat{\partial}f$ is the same as the usual convex subdifferential, i.e.,
^$\hat{\partial}f(x)=\partial f(x):=\{s\in\mathbb{R}^{n}\mid f(y)\geq f(x)+\mathopen{}\mathclose{{}\left\langle s,y-x}\right\rangle,\forall y\in\mathbb{R}^{n}\},\;\;\forall x\in\mathbb{R}^{n},$
and if $f$ is differentiable, Fréchet subdifferential reduces to the usual gradient, i.e., ^$\hat{\partial}f(x)=\{\nabla f(x)\},\forall x\in\mathbb{R}^{n}$.
Definition 2.1 (Slope [21, Section 2]).
@@
If $f$ is smooth, the slope $|\nabla f|(\bar{x})$ is simply the norm of its gradient $\|\nabla f(\bar{x})\|$. If $f(x)$ is convex, the slope $|\nabla f|(\bar{x})$ is equivalent to the length of the minimal norm element in the subdifferential, i.e., $|\nabla f|(\bar{x})=\mathrm{dist}(0,\partial f(\bar{x}))$. If it is $\rho$-weakly convex (i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex), the slope $|\nabla f|(\bar{x})$ is the same as the length of the minimal norm element in the Fréchet subdifferential, i.e., $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$ (see [38, 39] for more details). We summarize these useful properties into a lemma below; a brief proof is presented in Section B.2 for self-completeness:
@@
    If $f$ is differentiable, then $|\nabla f|(\bar{x})=\|\nabla f(\bar{x})\|$;
@@
We then introduce a key result from [40, Lemma 2.5] that will serve as a crucial step to estimate the distance to the optimal solution set.
Lemma 2.2 (Lemma 2.5 [40]).
@@
Then the sublevel set $[f\leq\alpha]$ is nonempty and $\mathrm{dist}(x,[f\leq\alpha])\leq(f(x)-\alpha)/r.$
@@
Finally, we review Ekeland’s variational principle [41].
Theorem 2.2 (Ekeland’s variational principle [41]).
@@
From the definition of slope (Definition 2.1), one immediate consequence of Theorem 2.2 is that the slope of $y$ is also bounded by $\rho$ as
@@ 3 Relationships between regularity conditions
In this section, we aim to show the equivalence between different regularity conditions in the class of weakly convex (potentially nonsmooth) functions. A function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is called $\rho$-weakly convex if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. The class of weakly convex functions is very rich: it includes all convex functions, $L$-smooth functions, certain compositions of convex functions with smooth functions, and many cost functions in modern machine learning applications; we refer interested readers to [27, 28] for more details.

### v3

@@ Abstract
keywords:
Error bound, Polyak-Łojasiewicz inequality, Quadratic growth, Proximal point method, Linear convergence, Weakly convex functions
@@ 1 Introduction
With the success of machine learning, (sub)gradient-type algorithms, such as (sub)gradient descent, coordinate descent, Frank-Wolf algorithm, and other variants, have gained much attention over the past decades [beck2017first, bottou2018optimization]. For smooth and/or convex cases, (sub)gradient methods are most well-understood. It is well-known that the basic gradient descent algorithm achieves linear convergence for minimizing smooth and strongly convex functions [nesterov2018lectures]. However, strong convexity is often not satisfied in practice, and many fundamental models in machine learning lack this good property [agarwal2010fast].
@@
This motivates the study of different regularity conditions that are weaker than strong convexity, such as Polyak-Łojasiewicz inequality [polyak1963gradient], restricted secant inequality [zhang2013gradient], subdifferential error bound [drusvyatskiy2018error], and quadratic growth [necoara2019linear]. These regularity conditions go beyond strong convexity, and they mainly involve the relationship concerning the distance to the solution set, the cost value gap, and the length of the (sub)gradient. They often play a crucial step in many algorithm analyses. It has been shown that under those regularity conditions, a range of first-order optimization algorithms still enjoy linear convergence in many cases; see e.g., [polyak1963gradient, karimi2016linear, cui2019r, drusvyatskiy2018error, diaz2023optimal, garber2023linear, liao2023overview, ding2023revisiting]. Moreover, these convergence analyses even become more elegant under suitable regularity conditions; examples include the gradient descent method for smooth optimization [nesterov2018lectures, Theorem 2.1.15]; the proximal gradient method for nonsmooth convex optimization [drusvyatskiy2018error, Theorem 3.2]; and the proximal point method (PPM) to solve $0\in T(x)$ where $T$ is a maximal monotone set value mapping [leventhal2009metric].
@@
Recent studies have revealed the relationship between some regularity conditions under various assumptions on the smoothness and convexity of the objective function [karimi2016linear, drusvyatskiy2018error, artacho2008characterization, bolte2017error, ye2021variational, zhu2023unified, drusvyatskiy2021nonsmooth, guille2021study, rebjock2023fast]. We postpone the comparison between different studies after we formally define regularity conditions (see Remark 1). Different conditions are useful to prove the convergence of various algorithms. For example, the subdifferential error bound yields a clean proof for the proximal point method (PPM) [rockafellar1976monotone, leventhal2009metric]; the step-length type error bound becomes useful for the proximal gradient method [drusvyatskiy2018error]; Polyak-Łojasiewicz inequality leads to a clean proof of the gradient descent [polyak1963gradient, karimi2016linear]; quadratic growth plays an important role in the analysis of proximal bundle methods [diaz2023optimal] and spectral bundle methods [ding2023revisiting, liao2023overview]. In some cases, it may be easier to establish one regularity condition than another one and then use the equivalence between the two to derive convergence results, e.g. [cui2016asymptotic, Proposition 3.3]. Therefore, understanding the relationship or equivalence between common regularity conditions becomes crucial in algorithm analyses.
@@
Despite the results mentioned above on the equivalence [karimi2016linear, drusvyatskiy2018error, artacho2008characterization, bolte2017error, ye2021variational, zhu2023unified, drusvyatskiy2021nonsmooth, guille2021study, rebjock2023fast], their scope remains limited to one of the following scenarios: 1) smooth functions, or 2) nonsmooth but convex functions. While smooth and convex (but not strongly convex) problems cover a variety of applications, practical applications routinely deal with problems that lack both qualities (e.g., training nonsmooth and nonconvex deep networks). Recent studies have further identified another amenable problem class: weakly convex functions [davis2019stochastic]. This class of functions includes all convex functions, $L$-smooth functions, certain compositions of convex functions with smooth functions, and many cost functions in modern machine learning [drusvyatskiy2020subgradient, atenas2023unified]. Some algorithms for convex functions can also be extended to solve weakly convex functions, e.g., the classical bundle method [liang2023proximal], the stochastic projected subgradient method [davis2019proximally], and the damped proximal augmented Lagrangian method [dahal2023damped]. A recent study also proposes a unified analysis framework to analyze the convergence of algorithms designed for weakly convex functions [atenas2023unified]. In this paper, we aim to generalize the previous results on the equivalence of different regularity conditions [karimi2016linear, drusvyatskiy2018error, artacho2008characterization, bolte2017error, ye2021variational, zhu2023unified, drusvyatskiy2021nonsmooth] from the class of smooth functions (and the class of convex functions) to the class of nonsmooth weakly convex functions. These findings will benefit algorithm analyses from various perspectives.
@@
To demonstrate the equivalence between different regularity conditions, we revisit the classical (inexact) PPM, a conceptually simple algorithm that serves as a guideline for many algorithm developments [drusvyatskiy2017proximal]. In particular, we provide a simple new proof for the linear convergence in both the convex and weakly convex settings using Polyak-Łojasiewicz inequality, subdifferential error bound, and quadratic growth. Our proof is partially inspired by a recent result on spectral bundle methods [ding2023revisiting, liao2023overview] and the analysis in [leventhal2009metric]. As a by-product of our investigation, we clarify some subtleties in inexact PPM in terms of the inexactness and feasibility, which are less emphasized in the literature. In this paper, our contributions are as follows.
@@
  • We first clarify the relationship among common regularity conditions, including strongly convex, restricted secant inequality, subdifferential error bound, Polyak-Łojasiewicz inequality, and quadratic growth, in the class of weakly convex functions (Theorem 3.1). Most of the proofs follow from some elementary analysis. The connection between Polyak-Łojasiewicz inequality and subdifferential error bound utilizes the notion of slope in [drusvyatskiy2021nonsmooth, Section 2] and requires some sophisticated arguments.We also derive the coefficient conversion among different conditions (see Table 1).
@@
  • Building upon the equivalence for different regularity conditions, we present a simple and modular proof for linear convergence of the PPM for convex optimization [rockafellar1976monotone] (Theorem 4.2). The linear convergence also extends to the class of weakly convex functions with a proper initialization (Theorem 4.3). Our linear convergence result allows non-unique minimizers which is more general than a classical assumption of local Lipschitz continuity in [rockafellar1976monotone] and also relaxes an upper Lipschitz condition in [luque1984asymptotic].
@@
  • Finally, we clarify some subtleties in the inexact PPM regarding inexactness stopping criteria and feasibility (Section 5.1). We further establish the linear convergence of inexact PPM (Theorem 5.3) under standard assumptions on the inexactness of the subproblem. Our proof is clean and modular, which combines a key inequality in [luque1984asymptotic] with Theorem 4.2.
@@
A preliminary summary of this work appeared in [pmlr-v242-liao24a]. In this paper, we 1) complete the detailed proofs for all theorems and derive coefficient conversion among different conditions in Table 1, 2) extend the linear convergence result of the PPM from optimization on convex functions to weakly convex functions (Theorem 4.3), and 3) provide extra technical discussions on inexact PPM (Section 5.1). The rest of this paper is structured as follows. Section 2 summarizes the linear convergence of gradient descent as motivation and introduces some preliminaries in nonsmooth optimization techniques. Section 3 presents the relationship among different regularity conditions. Sections 4 and 5 focus on the (inexact) PPM and establish the sublinear and linear convergences. Section 6 presents three numerical experiments. Section 7 concludes this paper. Some auxiliary proofs and discussions are provided in Appendices D, A, D, C and B.
@@ 2.1 Motivation: Linear convergence of gradient descent algorithms
where $t_{k}>0$ is the step size. A textbook result says that when choosing a constant step size $t_{k}=\frac{1}{L}$, the GD iterates converge to $f^{\star}$ with a sublinear rate (precisely, $f(x_{k})-f^{\star}\leq L\mathrm{dist}^{2}(x_{0},S)/(2k)$); see e.g., [nesterov2018lectures, Corollary 2.1.2]. If the function $f$ is strongly convex, GD achieves a global linear convergence [nesterov2018lectures, Theorem 2.1.15].
However, the assumption of strong convexity is often not satisfied in practice. It is known that some alternative weaker assumptions are sufficient for linear convergences. We here introduce two notions: restricted secant inequality (RSI) [zhang2013gradient], and Polyak-Łojasiewicz (PL) inequality [polyak1963gradient]. A differentiable function $f$ satisfies RSI if there exists $\mu_{\mathrm{r}}>0$ such that
@@
Note that both RSI 2 and PL 3 imply that any stationary point of $f$ is a global minimum. However, they do not imply the uniqueness of stationary points or the convexity of the function. One can think that RSI 2 (resp. PL 3) requires that the gradient $\nabla f(x)$ grows faster than that of quadratic functions when moving away from the solution set $S$ (resp. the optimal value $f^{\star}$). Linear convergence of GD under the PL inequality was first proved in [polyak1963gradient], and linear convergence under RSI was discussed in [guille2022gradient, Proposition 1] and [zhang2020new, Proposition 1]. We summarize a simple version below.
@@
Thanks to RSI 2 and PL inequality 3, the proof of Theorem 2.1 is very elegant, taking only a few lines. For self-completeness, we provide a simple proof and some additional discussions in Section B.1. In particular, the RSI 2 leads to a quick proof of 4a, and the PL 3 allows for a simple proof of 4b. It is known that for $L$-smooth convex functions, the two conditions RSI 2 and PL 3 are equivalent (cf. [karimi2016linear, Theorem 2]). Some recent studies, such as [bolte2017error, necoara2019linear, zhang2017restricted], have explored the relationships among different regularity conditions for linear convergences. A nice summary appeared in [karimi2016linear, Theorem 2], but it only works in the context of $L$-smooth functions.
@@ 2.2 The notion of slope and Ekeland’s variational principle
We first introduce the Fréchet subdifferential. Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed function. We define the Fréchet subdifferential (see e.g., [li2020understanding, Page 27]):
$\hat{\partial}f(x):=\mathopen{}\mathclose{{\left\{s\in\mathbb{R}^{n}\mid\liminf_{y\to x}\frac{f(y)-f(x)-\mathopen{}\mathclose{{\left\langle s,y-x}}\right\rangle}{\|y-x\|}\geq 0}}\right\}.$
If $f$ is convex, the Fréchet subdifferential $\hat{\partial}f$ is the same as the usual convex subdifferential, i.e.,
$\hat{\partial}f(x)=\partial f(x):=\{s\in\mathbb{R}^{n}\mid f(y)\geq f(x)+\mathopen{}\mathclose{{\left\langle s,y-x}}\right\rangle,\forall y\in\mathbb{R}^{n}\},\;\;\forall x\in\mathbb{R}^{n},$
and if $f$ is differentiable, Fréchet subdifferential reduces to the usual gradient, i.e., $\hat{\partial}f(x)=\{\nabla f(x)\},\forall x\in\mathbb{R}^{n}$.
Definition 1 (Slope [drusvyatskiy2021nonsmooth, Section 2]).
@@
If $f$ is smooth, the slope $|\nabla f|(\bar{x})$ is simply the norm of its gradient $\|\nabla f(\bar{x})\|$. If $f(x)$ is convex, the slope $|\nabla f|(\bar{x})$ is equivalent to the length of the minimal norm element in the subdifferential, i.e., $|\nabla f|(\bar{x})=\mathrm{dist}(0,\partial f(\bar{x}))$. If it is $\rho$-weakly convex (i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex), the slope $|\nabla f|(\bar{x})$ is the same as the length of the minimal norm element in the Fréchet subdifferential, i.e., $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$ (see [ioffe2017variational, drusvyatskiy2013slope] for more details). We summarize these useful properties into a lemma below; a brief proof is presented in Section B.2 for self-completeness:
@@
    If $f$ is smooth, then $|\nabla f|(\bar{x})=\|\nabla f(\bar{x})\|$;
@@
We then introduce a key result from [drusvyatskiy2015curves, Lemma 2.5] that will serve as a crucial step to estimate the distance to the optimal solution set.
Lemma 2.2 (Lemma 2.5 [drusvyatskiy2015curves]).
@@
Then the sublevel set $f\leq\alpha]$ is nonempty and $\mathrm{dist}(x,[f\leq\alpha])\leq(f(x)-\alpha)/r.$
@@
Finally, we review Ekeland’s variational principle [ekeland1974variational].
Theorem 2.2 (Ekeland’s variational principle [ekeland1974variational]).
@@
From the definition of slope (Definition 1), one immediate consequence of Theorem 2.2 is that the slope of $y$ is also bounded by $\rho$ as
@@ 3 Relationships between regularity conditions
In this section, we aim to show the equivalence between different regularity conditions in the class of weakly convex (potentially nonsmooth) functions. A function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is called $\rho$-weakly convex if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. The class of weakly convex functions is very rich: it includes all convex functions, $L$-smooth functions, certain compositions of convex functions with smooth functions, and many cost functions in modern machine learning applications; we refer interested readers to [drusvyatskiy2020subgradient, atenas2023unified] for more details.